### v2

# VoiceCloak: A Multi-Dimensional Defense Framework against Unauthorized Diffusion-based Voice Cloning

Abstract.
Diffusion Models (DMs) have achieved remarkable success in realistic voice cloning (VC), while they also increase the risk of malicious misuse. Existing proactive defenses designed for traditional VC models aim to disrupt the forgery process, but they have been proven incompatible with DMs due to the intricate generative mechanisms of diffusion. To bridge this gap, we introduce VoiceCloak, a multi-dimensional proactive defense framework with the goal of obfuscating speaker identity and degrading perceptual quality in potential unauthorized VC. To achieve these goals, we conduct a focused analysis to identify specific vulnerabilities within DMs, allowing VoiceCloak to disrupt the cloning process by introducing adversarial perturbations into the reference audio. Specifically, to obfuscate speaker identity, VoiceCloak first targets speaker identity by distorting representation learning embeddings to maximize identity variation, which is guided by auditory perception principles. Additionally, VoiceCloak disrupts crucial conditional guidance processes, particularly attention context, thereby preventing the alignment of vocal characteristics that are essential for achieving convincing cloning. Then, to address the second objective, VoiceCloak introduces score magnitude amplification to actively steer the reverse trajectory away from the generation of high-quality speech. Noise-guided semantic corruption is further employed to disrupt structural speech semantics captured by DMs, degrading output quality. Extensive experiments highlight VoiceCloak’s outstanding defense success rate against unauthorized diffusion-based voice cloning.
Digital forensics, Proactive defense, Diffusion models, Adversarial examples, Voice clone
ccs: Security and privacy Social aspects of security and privacyccs: Applied computing Sound and music computingcopyright: none
1. Introduction
Diffusion Models (DMs) (Ho et al., 2020; Song et al., 2020a, b; Rombach et al., 2022) have recently emerged as powerful generative tools, achieving unprecedented success within speech synthesis, particularly for realistic voice cloning (VC). Their iterative denoising process enables fine-grained control over the generation procedure, yielding synthetic speech with remarkable naturalness, detail, and fidelity to human voice (Popov et al., 2022; Kong et al., 2020; Jeong et al., 2021; Shen et al., 2024). However, the open-source availability and ease of use of these models intensify concerns about potential misuse. Attackers can synthesize highly realistic voice replicas from short public audio clips, as depicted in Figure 1(a), enabling sophisticated fraud and circumvention of voiceprint authentication.
Figure 1. Illustration of diffusion-based voice cloning malicious misuse. (a) Voice forgery enables threats of fraud. (b) Traditional methods struggle due to ineffective disruptive gradients. (c) Audio protected by VoiceCloak resists high-fidelity cloning.
To counter such unauthorized use, two main defense paradigms are introduced, including forgery detection and proactive disruption. Existing forgery detection methods for synthesized speech (Jung et al., 2022; Zhou and Lim, 2021; Wu et al., 2024) are reactive, often failing to prevent harm after dissemination. This highlights a critical need for proactive defenses that mitigate these security risks before they manifest. To this end, several studies (Huang et al., 2021; Yu et al., 2023; Chen et al., 2024; Li et al., 2023) have explored proactive disruption by introducing elaborately crafted, imperceptible, adversarial perturbations into reference audio. These perturbations aim to compromise the functionality of either the voice decoder or the speaker identity encoder, thereby disrupting the cloning process.
While effective on traditional conversion architectures, existing proactive defense methods fail to defend against diffusion-based voice cloning. Figure 1(b) illustrates that the incompatibility arises from gradient vanishing and dynamic conditioning, which present unique challenges for traditional defense methods. Specifically, (1) Established methods rely on computing gradients via a single forward pass of the generator, aligning perturbations with the generator’s complete structure. However, the DMs’ long, multi-step sampling process creates a deep computational graph prone to vanishing gradients. This makes single-pass gradient information ineffective at disrupting the full generation trajectory (Kang et al., 2023). (2) Other strategies target specific subnetworks, such as encoders designed to process inputs like identity or content independently locally. Conversely, DMs often employ dynamic conditioning mechanisms, where the focus can shift across conditional aspects. Consequently, localized methods designed to target static modules struggle to achieve global disruption in DMs. These fundamental incompatibilities underscore the need for novel strategies tailored to this generative paradigm.
In light of the incompatibility of prior defenses with the diffusion paradigm, we introduce VoiceCloak, a novel proactive defense framework designed to achieve two primary objectives against unauthorized Diffusion-based Voice Cloning (VC): Speaker Identity Obfuscation and Perceptual Fidelity Degradation. Driven by these goals, we conduct a targeted analysis to identify corresponding intrinsic vulnerabilities within the Diffusion Models (DMs). Based on this analysis, we design specific optimization objectives for the adversarial perturbation. The perturbation then exploits analysis results to effectively disrupt the diffusion-based voice cloning pipeline.
To achieve identity obfuscation, VoiceCloak first directly manipulates speaker identity representations within a universal embedding space to prevent DMs from extracting the target identity signature. We further optimize a loss term derived from psychoacoustic principles to maximize perceptual identity distance. Recognizing that convincing mimicry requires precise conditional guidance via attention mechanisms. This dependency makes attention a key for disrupting identity rendering. To exploit this, we introduce attention context divergence to prevent the attention mechanism from correctly utilizing the contextual information derived from the reference audio, thus disrupting the accurate alignment of speaker style with content and hindering convincing cloning.
Simultaneously, to achieve fidelity degradation, our analysis focused on vulnerabilities within the core generative process itself. We identified the precision of the iterative denoising trajectory as crucial for high-fidelity generation. Therefore, we employ score magnitude amplification (SMA) to exploit this sensitivity, actively steering the generation path away from high-quality speech regions. Additionally, acknowledging that the U-Net’s internal understanding of structural speech semantics contributes to naturalness, we utilize noise-guided semantic corruption to disrupt the capture of coherent structural features and promote incoherence within the noise semantic space, degrading the final audio quality. These goal-driven strategies, each originating from an analysis of vulnerabilities corresponding to our defense objectives, form the overall VoiceCloak framework. Extensive experiments confirm the superior defense efficacy of our method against diffusion-based voice cloning attacks under equivalent perturbation budgets.
Our contributions are summarized as follows:
  • We propose VoiceCloak, a novel defense framework against Diffusion-Based VC that prevents unauthorized voice ”theft” by exploiting intrinsic diffusion vulnerabilities through multi-dimensional adversarial interventions.
  • We introduce auditory-perception-guided adversarial perturbations into speaker identity representations and disrupt the diffusion conditional guidance process to effectively distort identity information in synthesized audio.
  • We present SMA, which controls the score function to divert the denoising trajectory, complemented by a semantic function designed to adversarially corrupt structural semantic features within the U-Net, degrading the perceptual fidelity of the forged audio.
2. Related Work
2.1. Audio Diffusion Models
Diffusion Models (Ruan et al., 2023; Kawar et al., 2023; Ruiz et al., 2023) a dominant force in generative modeling, demonstrating extraordinary performance across multi-modal tasks, significantly advancing audio synthesis tasks like text-to-speech (Popov et al., 2021; Jeong et al., 2021) and unconditional audio generation (Kong et al., 2020; Liu et al., 2023). Particularly within voice cloning (VC), diffusion-based methods(Popov et al., 2022; Choi et al., 2024), mostly leveraging score-based formulations via stochastic differential equations (SDEs) (Song et al., 2020b), now yield outputs with remarkable naturalness and speaker fidelity. While impressive, this state-of-the-art performance significantly heightens concerns regarding potential misuse, directly motivating proactive defense strategies such as the framework proposed herein.
2.2. Proactive Defense via Adversarial Examples
Beyond passive DeepFake detection, proactive defenses aim to preemptively disrupt malicious syntheses, such as Voice Cloning (VC) or Text-to-Speech (TTS), by introducing carefully crafted, often imperceptible adversarial perturbations to the original audio.
Initial efforts by Huang et al. (Huang et al., 2021) demonstrated the feasibility of perturbing reference audio to disrupt VC, though balancing effectiveness and imperceptibility remained a challenge. Subsequent research sought to improve this balance. For instance, VoiceGuard (Li et al., 2023) focused on imperceptibility, optimizing time-domain perturbations guided by psychoacoustic models to stay below auditory thresholds. AntiFake (Yu et al., 2023) pursued similar goals, incorporating a human-in-the-loop process to iteratively refine the trade-off between perceived identity protection and audio quality. Addressing efficiency, Dong et al. (Dong et al., 2024) proposed a GAN-based generator to create perturbations more rapidly than typical iterative optimization methods.
Despite these advancements, existing proactive strategies were largely developed for prior generative architectures (e.g., GANs). They often overlook the unique denoising processes and internal structures characteristic of Diffusion Models (DMs), which differ fundamentally from earlier paradigms. Recognizing this critical gap, we propose an effective defense specifically designed to counter the challenges posed by diffusion-based voice cloning.
3. Preliminaries
3.1. Score-based Diffusion
Unlike the original denoising diffusion model that utilizes a discrete-time diffusion process by Markov chains (Ho et al., 2020), the score-based generative models leverage stochastic differential equations (Song et al., 2020b) to define continuous-time diffusion processes. The forward process progressively adds noise to the original audio $x_{0}\sim p_{0}(x)$ to transform the complex data distribution of $p_{0}(x)$ into a simple prior Gaussian distribution. The evolution of noisy audio $x(t)$ is governed by an SDE:
(1) $d\mathbf{x}=f(\mathbf{x},t)dt+g(t)d\mathbf{w},$
where $f(\mathbf{x},t)$ is the drift coefficient, $g(t)$ is the diffusion coefficient controlling the noise level and $\mathbf{w}$ is a standard Wiener process. The corresponding reverse-time SDE that transforms $x_{T}$ back into $p_{0}(x)$ can be expressed as:
(2) $d\mathbf{x}=\left[f(\mathbf{x},t)-g(t)^{2}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})\right]dt+g(t)d\bar{\mathbf{w}},$
where $\bar{\mathbf{w}}$ is the reverse-time Wiener process and $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ represents the score-function. Then, the score-based diffusion model is trained to estimate the score function:
(3) $\theta^{*}=\arg\min_{\theta}\int_{0}^{1}\lambda_{t}\mathbb{E}_{\mathbf{x}_{t}}\left\|s_{\theta}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})\right\|_{2}^{2}dt,$
where the expectation is taken over the data distribution $p_{0}(\mathbf{x}_{0})$ and the transition kernel $p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$.
3.2. Adversarial Vulnerability Analysis
As mentioned before, the optimization of perturbation is guided by two core objectives: (1) Speaker Identity Obfuscation and (2) Perceptual Fidelity Degradation. The design of these perturbations stems from a targeted analysis to identify and exploit specific vulnerabilities within the Diffusion Model (DM) generative process itself.
A primary objective of VoiceCloak is speaker identity obfuscation. Convincing identity mimicry in Diffusion Models (DMs) necessitates precise conditional control, guided by fine-grained acoustic details perceived and modeled from the reference audio $x_{ref}$. This guidance critically relies on mechanisms that align the target speaker’s acoustic characteristics from $x_{ref}$ with the corresponding phonetic content from $x_{src}$. Specifically, the attention block within U-Net executes this alignment, establishing the crucial mapping between source content elements and relevant acoustic style features derived from $x_{ref}$. This content-to-style mapping creates a key vulnerability: inaccurate attention-based alignment directly compromises the successful rendering of the target speaker’s identity. Consequently, disrupting this pathway offers a strategy for identity obfuscation.
Complementary to identity obfuscation, the second objective is perceptual fidelity degradation, intended to diminish the usability of potentially synthesized speech. This involves targeting core mechanisms underpinning the high quality achieved by DMs. Central to this is the iterative denoising process. Diffusion-based VC synthesizes high-fidelity audio by progressively refining a sample, starting from noise $x_{T}\sim\mathcal{N}(0,I)$, guided at each timestep $t$ towards the real data distribution. Consequently, a well-trained model effectively learns a precise reverse denoising trajectory. Faithfully executing this learned trajectory is essential for generating the natural outputs. This reliance on trajectory precision presents an exploitable vulnerability. Therefore, adversarially diverting the trajectory can disrupt convergence towards the high-fidelity audio region on the manifold.
Additionally, the perceived quality and naturalness of synthesized speech are closely tied to the coherent processing of structural information within the U-Net architecture (Ronneberger et al., 2015). The U-Net generates hierarchical internal feature maps capturing essential data patterns. Prior attribute editing work (Oh et al., 2024; Tumanyan et al., 2023) confirms these internal features are controllable and encode crucial semantic and acoustic details impacting perceptual quality. Adversarially corrupting these representations can directly impair the model’s capacity to synthesize coherent and natural-sounding speech. Therefore, targeting these U-Net features offers a complementary strategy. This analysis reveals critical vulnerabilities in DMs, providing targeted avenues for adversarial intervention.
4. Methodology
Figure 2. Overview of the proposed VoiceCloak framework. Perturbation optimization is guided by gradients from $\mathcal{L}_{total}$, aggregating contributions from four core components targeting two defense objective: (1) Identity Obfuscation, achieved via Opposite-Gender Centroid Guidance and Attention Context Divergence; and (2) Perceptual Fidelity Degradation, implemented through Score Magnitude Amplification and Noise-Guided Semantic Corruption.
In this section, we elaborate on the design and implementation details of the proposed VoiceCloak framework. To provide the necessary background, the problem formulation is established in Section 4.1. As shown in Figure 2, the VoiceCloak consists of four sub-modules, which will be introduced separately next.
4.1. Problem Formulation
Assume that malicious users can obtain reference audio $x_{ref}$ of a target speaker from public platforms. Leveraging open-source Voice Cloning (VC) models, they can synthesize speech from $x_{src}$ that mimics the vocal characteristics of $x_{ref}$, denoted $\mathcal{VC}(x_{src},x_{ref},t)$. Our goal is to proactively safeguard $x_{ref}$ against unauthorized voice cloning. The core idea is to introduce a human-imperceptible perturbation $\delta$ into $x_{ref}$, generating an adversarial audio $x_{adv}=x_{ref}+\delta$, which disrupts the reverse diffusion process leading to corrupted conversions. Formally, our objective is to find an optimal perturbation $\delta$ that maximizes the dissimilarity between the outputs generated using $x_{ref}$ and $x_{adv}$:
(4) $\delta:=$ $\arg\max_{\delta}\ \mathcal{D}(\mathcal{VC}(x_{src},x_{ref},t),\ \mathcal{VC}(x_{src},x_{adv},t)),$
subject $\text{subject to}\quad\|\delta\|_{\infty}\leq\epsilon,$
where $\mathcal{D}(\cdot)$ is a function that measures the discrepancy, $t$ indicates the diffusion timestep, $\epsilon$ is the infinity norm bound budget to restrict $\delta$. Notably, we select DiffVC (Popov et al., 2022) as our primary threat model due to its representativeness in diffusion-based voice cloning. Our approach focuses on designing specific adversarial objectives that implicitly define $\mathcal{D}$ by exploiting the intrinsic vulnerabilities.
4.2. Adversarial Identity Obfuscation
Opposite-Gender Embedding Centroid Guidance.
Inspired by prior work showing dedicated speaker embeddings capture identity (Chen et al., 2024), directly manipulating these embeddings offers a direct approach, but its practical utility is hindered by poor transferability when attacking unknown or different encoder models.
Therefore, we explore leveraging speech Representation Learning (Chen et al., 2022b) to extract robust general acoustic representations that inherently capture speaker identity cues yet exhibit reduced downstream dependency. Specifically, we select WavLM (Chen et al., 2022a) as our representation extractor, denoted as $\mathcal{R}(\cdot)$. We then apply adversarial perturbations within this more general feature space to enhance defense transferability. As a baseline approach, we consider maximizing the distance between the perturbed audio $x_{adv}$ and the original one $x_{ref}$ using an untargeted objective:
(5) $\mathcal{L}_{ID}=1-SIM(\mathcal{R}(x_{adv}),\mathcal{R}(x_{ref})),$
where $SIM(x_{1},x_{2})$ represents the cosine similarity metric. Maximizing this effectively pushes $x_{adv}$ away from the reference identity representation domain. However, the untargeted objective lacks specific directionality in high-dimensional space. So we further incorporate a targeted component guided by psychoacoustic principles.
Psychoacoustic studies (Kreiman and Sidtis, 2011; Lavner et al., 2000; Reynolds et al., 2000) indicate that listeners are highly sensitive to specific acoustic cues for speaker identity, with significant perceptual differences often observed between genders (e.g., in F0 and formant structures). Leveraging this, we hypothesize that guiding the adversarial audio’s features towards characteristics of the opposite gender will likely create the strongest perceptual contrast, enhancing identity obfuscation.
Based on this insight, we propose an auditory-perception-guided adversarial perturbation. Specifically, we first randomly select a speaker from the dataset whose gender is opposite to that of $x_{ref}$. To establish a representative identity embedding, we compute the centroid $\mathcal{C}_{opp}$ by averaging the embeddings of all utterances in $\mathcal{X}$ within the WavLM feature space $\mathcal{R}$:
(6) $\mathcal{C}_{opp}=\frac{1}{N}\sum_{x_{i}\in\mathcal{X}}\mathcal{R}(x_{i}),$
(7) $\mathcal{L}_{ID}=1-SIM(\mathcal{R}(x_{adv}),\mathcal{R}(x_{ref}))+\underbrace{SIM(\mathcal{R}(x_{adv}),\mathcal{C}_{opp})}_{\text{Gender}},$
where $\mathcal{X}$ is the set of opposite gender utterances and $N$ represents length of $\mathcal{X}$.
Finally, minimizing the Eq. 7 directs the optimization process to simultaneously dissociate from the original speaker identity and converge towards a region selected based on psychoacoustic principles to maximize identity ambiguity.
Attention Context Divergence.
Building upon the motivation to disrupt conditional guidance via attention for identity obfuscation (Section 3.2), we introduce Attention Context Divergence. This strategy targets the attention mechanism, specifically interfering with how contextual information from the reference audio $x_{ref}$ is utilized.
In DiffVC, conditional information from $x_{ref}$ is integrated via Linear-attention layers (Katharopoulos et al., 2020). Within the U-Net, latent codes representing the content $x_{src}^{t}$ are linearly projected to $Q$ matrix, whereas condition latents $x_{ref}^{t}$ are projected to generate the $K$ and $V$ matrices. Linear-attention first computes a context matrix by aggregating values ($V$) weighted by keys ($K$):
(8) $\mathcal{H}_{ctx}(x_{ref})=Softmax(\phi^{(l,t)}(x_{ref}^{t})W_{K}^{l})(\phi^{(l,t)}(x_{ref}^{t})W_{V}^{l})^{T},$
where $\mathcal{H}_{ctx}$ represents the context hidden state, $\phi^{(l,t)}(\cdot)$ is the deep features of the $l^{th}$ block in U-Net at timestep $t$, and $W_{K}^{l},W_{V}^{l}$ are the projection matrices. Before this context interacts with queries ($Q$) to obtain the attention output: $\mathcal{A}^{l}=(\mathcal{H}_{ctx})^{T}(W_{Q}^{l}\phi^{(l,t)}(x_{src}^{t}))$. This context representation dynamically summarizes the aspects of the reference speaker’s style deemed most relevant for synthesizing the current output part based on content queries. By applying a softmax function, we obtain an explicit probability distribution over sequence positions which represents the distribution of the relevant information. Our strategy, therefore, is to maximize the Kullback-Leibler (KL) divergence (Kullback and Leibler, 1951) between the context distribution derived from the original reference and that derived from the adversarial reference:
(9) $\mathcal{L}_{ctx}$ $=D_{KL}(P_{ref}\parallel P_{adv}),$
$\text{where}\ P_{adv}$ $=Softmax(\mathcal{H}_{ctx}(x_{adv})),$
Maximizing this divergence forces the attention focus pattern to deviate from the original, thereby hindering accurate style alignment.
To enhance the impact on identity information, we focus the adversarial perturbation on specific layers within the U-Net’s downsampling path. The rationale is that these earlier layers primarily process coarser structural features and lower-frequency components (Wang et al., 2024) often strongly associated with speaker timbre and identity. Therefore, focusing our proposed loss $\mathcal{L}_{ctx}$ on the attention layers within the U-Net’s downsampling path, the calculation of $P_{adv}$ can be restated as:
(10) $P_{adv}=Softmax(\sum_{l}\mathcal{H}_{ctx}^{l}(x_{adv})),$
where the layer index $l$ iterates over the set Down ($l\in\text{Down}$), representing the U-Net’s downsampling blocks.
4.3. Perceptual Fidelity Degradation
Score Magnitude Amplification.
To degrade perceptual fidelity by exploiting the sensitivity of the denoising trajectory, we introduce Score Magnitude Amplification. This design directly interferes with the score function $s_{\theta}$, which is estimated by the U-Net according to Eq. 3 and provides the essential drift term for the reverse SDE. We posit that the magnitude of $s_{\theta}$ relates to the strength of the drift guiding the sample toward the target data manifold. Exploiting this connection, the SMA objective involves maximizing the magnitude of the score prediction $s_{\theta}$:
(11) $\mathcal{L}_{score}=\mathbb{E}_{x_{src}^{t}\sim q(x_{src}^{t}|x_{src}^{0}),t\sim\mathcal{U}(0,T_{adv}]}[\|s_{\theta}(x_{src}^{t},x_{adv}^{t},t)\|_{2}],$
where $q(x_{src}^{t}|x_{src}^{0})$ is forward diffusion distribution, $\mathbb{E}[\cdot]$ calculates the average value, $T_{adv}$ stands for adversarial timesteps which will be discussed in Section 4.4. Optimizing the above formula introduces an erroneous drift strength. Consequently, the denoising trajectory is forcefully steered away from the intended high-quality manifold, resulting in a collapse in the perceptual quality.
Furthermore, the iterative nature of the diffusion process may amplify the impact of this manipulation. Perturbations introduced at earlier timesteps can propagate through subsequent steps. This cumulative effect thus enhances the efficacy of our adversarial strategy.
Noise-Guided Semantic Corruption.
Following the motivation outlined in Section 3.2, we introduce a bidirectional semantic interference strategy. The core idea is twofold: (1) compel the features generated with $x_{adv}$ to diverge from those generated using the original reference $x_{ref}$, and (2) concurrently guide these adversarial features towards a ”semantic-free” target state.
Specifically, consider a layer $l$ within the frozen U-Net $\mathcal{U}_{\theta}$ and timestep $t$, we extract the original features $f^{(l,t)}=\mathcal{U}_{\theta}^{l}(x_{src},x_{ref},t)$ conditioned on $x_{ref}$, and $f^{(l,t)}_{adv}=\mathcal{U}_{\theta}^{l}(x_{src},x_{adv},t)$ corresponding to the adversarial one. Furthermore, to define a ”semantic-free” target, we leverage the U-Net’s activation modes to unstructured information. We extract features $f^{(l,t)}_{noise}=\mathcal{U}_{\theta}^{l}(x_{noise},x_{noise},t)$ by feeding Gaussian white noise $x_{noise}$ as both the source content and the reference condition. $f^{(l,t)}_{noise}$ can be considered to represent unstructured features and lack semantic information. The bidirectional objective aims to maximize the distance between $f^{(l,t)}_{adv}$ and $f^{(l,t)}$ while minimizing it with $f^{(l,t)}_{noise}$. This objective encourages the adversarial features to abandon the original semantic structure and move towards a state of incoherence. The objective is formalized as:
(12) $\mathcal{L}_{sem}=1-\cos(f^{(l,t)}_{adv},f^{(l,t)})+\underbrace{\cos(f^{(l,t)}_{adv},f^{(l,t)}_{noise})}_{\text{Sem-free}},$
where we employ the cosine distance metric $\cos(\cdot)$ as it emphasizes the structural similarity between high-dimensional features rather than their absolute error. For enhanced impact on perceptual quality, the $\mathcal{L}_{sem}$ is strategically applied to layers within the U-Net’s upsampling path. These layers are critical for reconstructing the fine-grained acoustic details that govern output naturalness and perceived quality.
4.4. Joint Optimization of Defense Objectives
The final adversarial perturbation $\delta$ for VoiceCloak is optimized to simultaneously achieve identity obfuscation (Section 4.2) and perceptual fidelity degradation (Section 4.3). The joint objectives of this comprehensive defense are formalized as follows
(13) $\mathcal{L}_{total}=$ $(\mathcal{L}_{ID},\mathcal{L}_{ctx},\mathcal{L}_{score},\mathcal{L}_{sem})\Lambda^{T},$
$\delta:=$ $\ \arg\max_{\delta}\ \mathcal{L}_{total},$
where $\Lambda=(\lambda_{ID},\lambda_{ctx},\lambda_{score},\lambda_{sem})$ controls the weight factors that balance the relative importance of these defenses.
The efficacy of our perturbation optimization is influenced by the choice of diffusion timesteps $T_{adv}$ used for gradient computation. Informed by prior work (Yu et al., 2024) indicating early denoising steps primarily reconstruct low-frequency signal structure, we concentrate our optimization on these initial steps to maximize the disruption of fundamental integrity and reduce computational overhead.
Table 1. Comparison of defense effectiveness and adversarial imperceptibility with SOTA methods. Higher values are better for metrics marked with $\uparrow$, and vice versa for those marked with $\downarrow$. The best result is marked in BOLD.
Datasets Methods Defense Effectiveness Imperceptibility
[…]
LibriTTS Undefended - 76.49% - 3.96 - - - -
Random Noise $l_{\infty}$ 2.01 55.20% 0.31 3.72 16.00% 3.37 1.35 34.80
Attack-VC 2.29 36.20% 0.31 3.57 30.40% 2.31 3.71 5.29
VoicePrivacy 2.26 20.80% 0.30 3.60 26.80% 2.99 1.37 33.25
VoiceGuard 2.08 16.49% 0.29 3.63 43.45% 2.15 4.39 10.58
Ours 2.12 11.40% 0.27 2.36 71.40% 3.22 1.29 33.53
VCTK Undefended - 63.68% - 3.41 - - - -
Random Noise $l_{\infty}$ 1.68 58.00% 0.35 3.16 11.38% 3.25 1.38 34.10
Attack-VC 2.05 38.50% 0.33 2.82 26.20% 2.25 3.80 4.50
VoicePrivacy 1.88 30.28% 0.35 2.77 39.06% 2.87 1.53 31.87
VoiceGuard 1.87 31.42% 0.32 3.02 22.11% 2.05 4.57 10.00
Ours 1.93 19.74% 0.29 2.51 63.41% 3.09 1.33 32.41
5. Experiments
5.1. Experimental Setup
Datasets. In our experiments, we utilize two primary English speech datasets, including LibriTTS (Zen et al., 2019) and VCTK corpus (Yamagishi et al., 2019). To generate adversarial references ($x_{adv}$), we select subsets of 479 utterances from the LibriTTS dev-clean subset and 500 utterances from VCTK, ensuring approximate gender balance ($50\%$ male/female). Additionally, one male and one female speaker are selected to provide the source content. The evaluation focused on gender-matched conversions.
Baseline Methods We compare VoiceCloak against existing voice protection methods. Given the lack of protection methods for diffusion-based voice conversion, we adopt the following methods for fair comparison: Attack-VC (Huang et al., 2021), VoicePrivacy (Chen et al., 2024), and VoiceGuard (Li et al., 2023). We also include a naive baseline: adding random Gaussian noise to $x_{ref}$ with an $l_{\infty}$-norm matched to our adversarial perturbations.
Evaluation Metrics We evaluate defense performance using objective metrics and a subjective user study (Sec. 5.2.3). Let $y$ and $y_{adv}$ be the audio synthesized using the original reference $x_{ref}$ and the defended $x_{adv}$, respectively. We use Automatic Speaker Verification (ASV) to quantify identity protection. We measure identity similarity by comparing the embedding of the defended output $y_{adv}$ against the target speaker’s average embedding. The acceptance rate of ASV is then calculated as the proportion of $y_{adv}$ samples whose score $>\tau_{ASV}$. To assess the overall Defense Success Rate (DSR), reflecting both identity obfuscation and quality degradation, we combine the ASV results with a naturalness score predictor NISQA (Mittag and Möller, 2021): successful defense requires failing ASV while also exhibiting low quality ($\mathrm{DS}=s_{ASV}<\tau_{ASV}\land\mathrm{NISQA}(y_{adv})<\tau_{q}$ where we set $\tau_{ASV}=0.25$ and $\tau_{q}=3.0$) Additional objective metrics include Dynamic Time Warping (DTW) (Sakoe and Chiba, 1978) and Structural Similarity Index Measure (SSIM) between $y$ and $y_{adv}$ spectrograms. Perturbation imperceptibility on $x_{adv}$ versus $x_{ref}$ is measured using Perceptual Evaluation of Speech Quality (PESQ) (Rix et al., 2001), Mel-Cepstral Distortion (MCD) (Kubichek, 1993) and Signal-to-Noise Ratio (SNR).
Implementation Details All audio data are uniformly resampled to 22050 Hz. We set the number of optimization iterations to 50 and a step size $\alpha=4\times 10^{-5}$. And $\delta$ is constrained within an $l_{\infty}$-norm ball of $\epsilon=0.002$. We set the adversarial and inference timesteps respectively $T_{adv}=6$ and $T=100$. The loss function combined identity and quality objectives with weights $\Lambda=(1.0,4.5,10,0.85)$. All experiments are conducted on NVIDIA RTX 3090 GPU with a fixed random seed.
5.2. Comparison and Analysis
5.2.1. Comparison with Baselines
Figure 3. Mel spectrograms with $F_{0}$ pitch contours (green lines), and inferred intonation of the corresponding words. Arrows indicate perceived intonation shifts. (Intonation aligns with the ground truth, which is marked by green arrows, and diverges, which is marked by red arrows.)
We conduct a quantitative comparison of our VoiceCloak against baseline methods on the LibriTTS and VCTK datasets, evaluating both defense effectiveness and perturbation imperceptibility. The results are shown in Table 1.
Defense Effectiveness. Our framework demonstrates exceptional efficacy in neutralizing the threat of diffusion-based voice cloning. This is most strikingly evidenced by the Defense Success Rate (DSR)-over 71% on LibriTTS and 63% on VCTK, significantly surpassing all baseline approaches. This success stems from our multi-faceted strategy targeting inherent vulnerabilities within the diffusion process. Specifically, the framework achieves significant identity obfuscation, drastically reducing speaker similarity with ASV down to $11\%$, indicating the failure of voice mimicry. Simultaneously, it ensures substantial perceptual quality degradation, reflected in low NISQA scores, a range typically considered unacceptable generation quality. This dual ability to effectively cripple both identity mimicry and audio usability distinguishes VoiceCloak from defenses that may struggle with diffusion models or focus primarily on one objective. The ability to cripple both identity mimicry and audio usability simultaneously underscores the benefit of our comprehensive design. As expected, the naive Random Noise baseline confirms that unstructured noise provides very limited protection.
Figure 3 further visualizes some disruption results with fundamental frequency ($F_{0}$) and intonations. Applying VoiceCloak protection results in forged speech markedly different from the Undefended forgery. The $F_{0}$ curve is notably degraded, appearing blurred and unpredictable, accompanied by highly inconsistent intonation changes.
Imperceptibility. Regarding perturbation imperceptibility, VoiceCloak achieves performance comparable to most baseline methods, as indicated by similar PESQ, SNR scores, and lowest spectral distortion (MCD). This contrasts with methods like VoiceGuard, which significantly sacrifices audio fidelity without commensurate defense gains. This underscores that directly targeting intrinsic vulnerabilities, rather than simply increasing perturbation magnitude, is a more effective strategy for defending diffusion models while maintaining reasonable imperceptibility.
5.2.2. Protecting Commercial Systems
Figure 4. Protecting commercial speaker verification APIs (Iflytek and Azure) from voice spoofing attacks: average similarity confidence scores (lower scores indicate better defense).
We evaluate our proposed method’s effectiveness in protecting commercial speaker verification (SV) systems, such as Iflytek and Azure platforms, against voice spoofing attacks. These APIs output a confidence score indicating speaker similarity, which higher scores increase the vulnerability to false acceptance from spoofing attacks. Therefore, successful protection aims to minimize this score when verifying protected audio against the original speaker’s enrollment. As shown in Figure 4 our proposed method consistently achieves the lowest confidence scores across both platforms. This indicates that VoiceCloak more effectively decouples the protected audio from the original speaker’s identity, underscoring its superior performance in practical, black-box scenarios.
5.2.3. User Study
Figure 5. User perceptual study results. (a) Timbre Dissimilarity Preference. (b) Corresponding results for perceived Naturalness Disruption.
To assess the human perceptual impact of our approach, we conduct a subjective study with 50 participants evaluating samples from VoiceCloak and baseline methods. Aggregated preferences are shown in Figure 5. The user study results demonstrate the perceptual impact of VoiceCloak. Participants consistently rated audio protected by our method as less similar in timbre to the target speaker and significantly less natural compared to baseline outputs, confirming the framework’s effectiveness in disrupting both identity and quality according to human listeners.
Table 2. Ablation study on the contribution of different settings for Identity Obfuscation. Checkmark ($\checkmark$) indicates the setting is used, while $\downarrow$ signifies that a lower value is more desirable. ”w/o” denotes the exclusion of the specified loss term.
Settings Defense Effectiveness Imperceptibility
[…]
- - 1.96 46.82% 0.30 3.66 22.58% 3.34 1.46 34.94
$\checkmark$ - 2.16 8.57% 0.30 3.57 27.74% 3.09 1.45 33.43
w/o Gender - 2.25 19.92% 0.30 3.60 14.40% 3.00 1.36 33.26
- $\checkmark$ 2.31 16.20% 0.27 2.96 62.57% 3.22 1.28 33.58
$\checkmark$ $\checkmark$ 2.13 11.00% 0.27 2.85 69.20% 3.21 1.29 33.48
5.3. Ablation Study
5.3.1. Effectiveness of Adversarial Identity Obfuscation
Table 3. Ablation study on the contribution of different settings for Perceptual Fidelity Degradation. Checkmark ($\checkmark$) indicates the setting is used, while $\downarrow$ signifies that a lower value is more desirable. ”w/o” denotes the exclusion of the specified loss term.
Settings Defense Effectiveness Imperceptibility
[…]
- - 1.99 45.00% 0.31 3.09 20.20% 3.38 1.35 34.82
$\checkmark$ - 2.42 31.80% 0.29 2.68 41.20% 3.30 1.30 34.03
- $\checkmark$ 2.23 23.00% 0.27 2.44 60.60% 3.28 1.29 33.83
- w/o Sem-free 2.28 26.36% 0.29 3.30 26.80% 3.00 1.36 33.28
$\checkmark$ $\checkmark$ 2.22 23.60% 0.27 2.10 57.80% 3.29 1.29 33.91
We first investigate the individual contributions of the loss components designed for Adversarial Identity Obfuscation (Section 4.2), namely $\mathcal{L}_{ID}$ and $\mathcal{L}_{ctx}$. The results are presented in Table 2. Employing $\mathcal{L}_{ID}$ markedly reduces the speaker verification accept rate. This confirms that directly manipulating the representation learning embedding space effectively disrupts recognizable identity features. Furthermore, the higher ASV (19.92%) when removing opposite-gender guidance indicates our psychoacoustically-motivated strategy provides directional control to specifically disrupt structural identity cues. The context divergence loss $\mathcal{L}_{ctx}$ also significantly contributes to defense, lowering both ASV and NISQA. This suggests that $\mathcal{L}_{ctx}$ interferes with the attention mechanism’s conditioning on the reference speaker’s style, leading to a boost in the overall DSR.
5.3.2. Effectiveness of Perceptual Fidelity Degradation
Table 3 shows the ablation study evaluating components designed for perceptual quality degradation. We observe that applying $\mathcal{L}_{score}$ alone forces the denoising trajectory away from the high-fidelity manifold via its amplified drift term. Consequently, NISQA drops to 2.68 and DTW distance increases. The semantic corruption loss $\mathcal{L}_{sem}$ demonstrates an even stronger individual impact on perceived quality. Its effectiveness stems from directly corrupting internal U-Net features, thereby impairing the model’s reconstruction of coherent, natural-sounding speech details. Critically, the necessity of guiding semantic features towards an incoherent state is evident from the ”w/o Sem-free” variant. Removing this noise-guidance component results in audio quality comparable to the baseline.
5.3.3. Perturbation Budget
Table 4. Performance of VoiceCloak under different perturbation budgets. ”Undefended” denotes voice conversion without VoiceCloak.
Budgets ($\epsilon$) Defense Effectiveness Imperceptibility
[…]
Undefended - 76.49% - 3.96 - - - -
0.0005 2.01 22.00% 0.29 2.86 41.20% 4.25 0.48 45.03
0.0010 2.05 16.60% 0.28 2.65 57.00% 3.81 0.79 39.26
0.0020 2.12 11.40% 0.27 2.36 71.20% 3.21 1.29 33.53
0.0050 2.23 6.20% 0.25 1.95 84.00% 2.34 2.34 25.94
0.0100 2.32 1.20% 0.24 1.75 90.80% 1.74 3.40 20.18
We investigated the influence of the perturbation budget $\epsilon$ on both defense effectiveness and adversarial imperceptibility, with results summarized in Table 4. The findings reveal a clear trade-off: increasing e directly enhances defense effectiveness but conversely degrades audio imperceptibility. The excessive distortion compromises the usability of $x_{adv}$ for legitimate purposes. We determined $\epsilon=0.0020$ to be a suitable default budget, providing strong protection while preserving acceptable adversarial quality.
5.3.4. Impact of the Sampling Step
Table 5. Ablation study on the effect of different diffusion inference steps $T$. ”Avg. RT” indicates the average runtime per generated sample.
Infer. Steps $T$ Defense Effectiveness Avg. RT (s)
[…]
6 1.72 11.80% 1.83 84.20% 0.32
18 1.94 10.40% 2.08 81.80% 0.71
30 2.03 11.40% 2.13 78.80% 1.09
100 2.13 11.00% 2.37 71.80% 3.28
200 2.15 13.40% 2.37 69.20% 6.57
Table 5 investigates the impact of varying inference steps $T$, addressing the realistic scenario where an attacker’s generation timestep is unknown. The results demonstrate strong stability in defense effectiveness regardless of the timesteps. Notably, the DTW score, tends to increase with more inference steps. We attribute this to the cumulative amplification effect discussed in Section 4.3, whereby induced trajectory deviations compound over longer inference paths.
6. Conclusion
In this paper, we introduced VoiceCloak, a novel, multi-dimensional proactive defense framework designed to safeguard against unauthorized diffusion-based voice cloning. We established that a comprehensive defense should achieve the dual objectives of speaker identity obfuscation and perceptual quality degradation by exploiting targeted intrinsic vulnerabilities. To achieve identity obfuscation, it directly manipulates speaker representations and interferes with the attention mechanism’s understanding of contextual information via attention context divergence. For perceptual quality degradation, it employs Score Magnitude Amplification to steer the reverse denoising trajectory and utilizes noise-guided semantic corruption to promote semantic incoherence by disrupting internal feature processing. Experiments demonstrate the effective defense performance of VoiceCloak, thereby mitigating the potential harms associated with high-quality diffusion-based voice mimicry.
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (No.62441237, No.62261160653, No.U2001202, No.62072480), the Guangdong Provincial Key Laboratory of Information Security Technology (No. 2023B1212060026).