### v4

# VoiceCloak: A Multi-Dimensional Defense Framework against Unauthorized Diffusion-based Voice Cloning

Abstract
Diffusion Models (DMs) have achieved remarkable success in realistic voice cloning (VC), while they also increase the risk of malicious misuse. Existing proactive defenses designed for traditional VC models aim to disrupt the forgery process, but they have been proven incompatible with DMs due to the intricate generative mechanisms of diffusion. To bridge this gap, we introduce VoiceCloak, a multi-dimensional proactive defense framework with the goal of obfuscating speaker identity and degrading perceptual quality in potential unauthorized VC. To achieve these goals, we conduct a focused analysis to identify specific vulnerabilities within DMs, allowing VoiceCloak to disrupt the cloning process by introducing adversarial perturbations into the reference audio. Specifically, to obfuscate speaker identity, VoiceCloak first targets speaker identity by distorting representation learning embeddings to maximize identity variation, which is guided by auditory perception principles. Additionally, VoiceCloak disrupts crucial conditional guidance processes, particularly attention context, thereby preventing the alignment of vocal characteristics that are essential for achieving convincing cloning. Then, to address the second objective, VoiceCloak introduces score magnitude amplification to actively steer the reverse trajectory away from the generation of high-quality speech. Noise-guided semantic corruption is further employed to disrupt structural speech semantics captured by DMs, degrading output quality. Extensive experiments highlight VoiceCloak’s outstanding defense success rate against unauthorized diffusion-based voice cloning. Audio samples of VoiceCloak are available at https://voice-cloak.github.io/VoiceCloak/.
1 Introduction
Diffusion Models (DMs) (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020; Rombach et al. 2022) have recently emerged as powerful generative tools, achieving unprecedented success within realistic voice cloning (VC). Their iterative denoising process enables generating speech with remarkable naturalness, detail, and fidelity to human voice (Popov et al. 2022; Kong et al. 2020; Jeong et al. 2021; Shen et al. 2024). However, the open-source availability and ease of use of these models intensify concerns about potential misuse. Attackers can synthesize highly realistic voice replicas from short public audio clips, as depicted in Figure 1(a), enabling sophisticated fraud and circumvention of voiceprint authentication.
Figure 1: Illustration of diffusion-based voice cloning malicious misuse. (a) Voice forgery enables threats of fraud. (b) Traditional methods struggle due to ineffective disruptive gradients. (c) Audio protected by VoiceCloak resists high-fidelity cloning.
To counter such unauthorized use, two main defense paradigms are introduced, including forgery detection and proactive disruption. Reactive detection methods (Jung et al. 2022; Zhou and Lim 2021; Wu et al. 2024) identify forgeries after they are crafted, often too late to prevent harm. This highlights the need for proactive defenses that disrupt the synthesis process itself. Prior proactive work (Huang et al. 2021; Yu, Zhai, and Zhang 2023; Chen et al. 2024; Li et al. 2023) has focused on adding imperceptible adversarial perturbations to reference audio by compromising the functionality of either the voice decoder or the speaker identity encoder.
However, existing defenses designed for prior architectures are largely ineffective against Diffusion Models (DMs). This incompatibility arises from two fundamental challenges: gradient vanishing and dynamic conditioning (Figure 1). Specifically, (1) the single forward pass gradient computation relied upon by many defenses become unreliable or vanish within the multi-step denoising process of DMs and the corresponding deep computational graph, rendering such single-pass gradient information ineffective for disrupting the full generation trajectory (Kang, Song, and Li 2023). (2) Strategies targeting specific subnetworks (e.g. speaker or content encoders) fail because DMs often employ dynamic conditioning mechanisms, which means no single modules solely responsible for condition processing. Consequently, methods targeting individual components struggle to cause global disruption. These fundamental incompatibilities underscore the need for novel strategies tailored to this generative paradigm.
In light of the incompatibility of prior defenses with the diffusion paradigm, we introduce VoiceCloak, a novel proactive defense framework designed for two primary objectives against unauthorized voice cloning: Speaker Identity Obfuscation and Perceptual Fidelity Degradation. Driven by these two objectives, we conduct an analysis to identify and exploit corresponding intrinsic vulnerabilities within Diffusion Models (DMs). Based on this analysis, we design specific optimization objectives for the protective perturbation to effectively disrupt the synthesis process.
To achieve identity obfuscation, VoiceCloak first directly manipulates speaker representations within a universal embedding space, guided by psychoacoustic principles to maximize perceived identity distance and hinder the DM’s identity signature extraction. Second, recognizing that convincing mimicry depends on attention mechanisms in conditional guidance to align speaker style with content, we exploit this by introducing attention context divergence. This design prevents the attention mechanism from correctly utilizing contextual information, thereby disrupting the alignment required for accurate cloning.
Simultaneously, to achieve fidelity degradation, we focus on vulnerabilities within the core generative process itself. First, we employ Score Magnitude Amplification (SMA) to exploit the sensitivity of the iterative denoising trajectory which is crucial for realistic output, steering the generation path away from high-fidelity regions. Furthermore, acknowledging that the U-Net’s understanding of high-level semantics governs output naturalness, we utilize noise-guided semantic corruption to disrupt the capture of structural features to promote incoherence within the noise semantic space and degrade generation quality. These goal-driven strategies which originates from adversarial analysis form a comprehensive defense. Extensive experiments confirm the superior defense efficacy against diffusion-based VC attacks under equivalent perturbation budgets.
Our contributions are summarized as follows:
  • We propose VoiceCloak, a novel defense framework against Diffusion-Based VC that prevents unauthorized voice ”theft” by exploiting intrinsic diffusion vulnerabilities through multi-dimensional adversarial interventions.
  • We introduce auditory-perception-guided adversarial perturbations into speaker identity representations and disrupt the diffusion conditional guidance process to effectively distort identity information in synthesized audio.
  • We present SMA, which controls the score function to divert the denoising trajectory, complemented by a semantic function designed to adversarially corrupt structural semantic features within the U-Net, degrading the perceptual fidelity of the forged audio.
2 Related Work
2.1 Audio Diffusion Models
Diffusion Models (Ruan et al. 2023; Kawar et al. 2023; Ruiz et al. 2023) a dominant force in generative modeling, demonstrating extraordinary performance across multi-modal tasks, significantly advancing audio synthesis tasks like text-to-speech (Popov et al. 2021; Jeong et al. 2021) and unconditional audio generation (Kong et al. 2020; Liu et al. 2023). Particularly within voice cloning (VC), diffusion-based methods(Popov et al. 2022; Choi, Lee, and Lee 2024), mostly leveraging score-based formulations via stochastic differential equations (SDEs) (Song et al. 2020), now yield outputs with remarkable naturalness and speaker fidelity. While impressive, this state-of-the-art performance significantly heightens concerns regarding potential misuse, directly motivating proactive defense strategies such as the framework proposed herein.
2.2 Proactive Defense via Adversarial Examples
Beyond passive DeepFake detection, proactive defenses aim to preemptively disrupt malicious syntheses, by introducing adversarial perturbations to the original audio. Early work (Huang et al. 2021) demonstrated the feasibility of this approach but struggled to balance effectiveness with imperceptibility. Subsequent research focused on improving this trade-off. Strategies included using psychoacoustic models to enhance imperceptibility (Li et al. 2023), incorporating human-in-the-loop refinement for better balance (Yu, Zhai, and Zhang 2023), , and employing GAN-based generators to improve efficiency (Dong et al. 2024).
Despite these advancements, prior proactive strategies were largely designed for earlier generative architectures like GANs. They often overlook the unique mechanisms and internal structures of Diffusion Models (DMs), limiting their applicability. Recognizing this critical gap, our work proposes a defense specifically tailored to the challenges of diffusion-based voice cloning.
3 Preliminaries
3.1 Score-based Diffusion
Score-based generative models define a continuous-time diffusion process using stochastic differential equations (SDEs) (Song et al. 2020). The forward process gradually perturbs clean data $x_{0}\sim p_{0}(x)$ into noise via the SDE:
$d\mathbf{x}=f(\mathbf{x},t)dt+g(t)d\mathbf{w},$ (1)
where $f(\mathbf{x},t)$ and $g(t)$ are the drift and diffusion coefficients, and $\mathbf{w}$ is a Wiener process. The corresponding reverse-time SDE that transforms $x_{T}$ back into $p_{0}(x)$ can be expressed as:
$d\mathbf{x}=\left[f(\mathbf{x},t)-g(t)^{2}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})\right]dt+g(t)d\bar{\mathbf{w}},$ (2)
where $\bar{\mathbf{w}}$ is the reverse-time Wiener process and $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ represents the score-function. Then, the score-based diffusion model is trained to estimate the score function:
$\arg\min_{\theta}\lambda_{t}\mathbb{E}_{p_{t}(x)}\left\|s_{\theta}(\mathbf{x}_{t},t)-\nabla\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})\right\|^{2}_{2},$ (3)
where the expectation is taken over the data distribution $p_{0}(\mathbf{x}_{0})$ and the transition kernel $p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$.
3.2 Adversarial Vulnerability Analysis
As mentioned before, the optimization of perturbation is guided by two core objectives: (1) Speaker Identity Obfuscation and (2) Perceptual Fidelity Degradation. The design of these perturbations stems from a targeted analysis to identify and exploit specific vulnerabilities within the Diffusion Model (DM) generative process itself.
A primary objective of VoiceCloak is speaker identity obfuscation. Convincing identity mimicry in Diffusion Models (DMs) depends on precise conditional control which is guided by acoustic details modeled from the reference audio $x_{ref}$. This guidance critically relies on mechanisms that align the target speaker’s acoustic characteristics from $x_{ref}$ with the phonetic content from $x_{src}$. Specifically, the attention block is responsible for executing this alignment. We identify this crucial acoustic style-to-content mapping as a key vulnerability,as inaccurate alignment directly compromises the successful rendering of the target speaker’s identity. Consequently, disrupting this attention-driven pathway offers a strategy for identity obfuscation.
Complementary to identity obfuscation, the second objective is perceptual fidelity degradation, aimed at diminishing the usability of any synthesized speech. This involves targeting the core denoising process of DMs. The high fidelity of Diffusion-based VC relies on the model learning a precise reverse denoising trajectory to progressively refine noisy sample $x_{t}$ towards the natural audio distribution. This reliance on trajectory precision presents an exploitable vulnerability. Therefore, adversarially diverting the trajectory can disrupt convergence towards the high quality audio region on the manifold.
Additionally, we target the U-Net’s internal feature representations as a vulnerability for degrading perceptual quality. Prior attribute editing work (Ronneberger, Fischer, and Brox 2015; Oh, Lee, and Lee 2024; Tumanyan et al. 2023) confirms that these hierarchical features within the U-Net are controllable and also encode crucial semantic and acoustic details that govern the coherence and naturalness of the synthesized speech. Therefore, adversarially corrupting these representations directly impair the model’s ability to synthesize natural-sounding speech, offering a complementary strategy for fidelity degradation. This analysis reveals critical vulnerabilities in DMs, providing targeted avenues for adversarial intervention.
4 Methodology
Figure 2: Overview of the proposed framework. Perturbation optimization is guided by gradients from $\mathcal{L}_{total}$, aggregating four targeting two objectives: (1) Identity Obfuscation (via Opposite-Gender Centroid Guidance and Attention Context Divergence) and (2) Perceptual Fidelity Degradation (Score Magnitude Amplification and Noise-Guided Semantic Corruption).
To provide the necessary background, the problem formulation is firstly established in Section 4.1. As shown in Figure 2, the VoiceCloak consists of four sub-modules, which will be introduced separately next.
4.1 Problem Formulation
Assume that malicious users can obtain reference audio $x_{ref}$ of a target speaker. Leveraging open-source voice cloning models, they can synthesize speech which mimics the vocal characteristics of $x_{ref}$, denoted $\mathcal{VC}(x_{src},x_{ref},t)$. Our goal is to proactively safeguard $x_{ref}$ against unauthorized voice cloning. We achieve this by introducing an imperceptible adversarial perturbation $\delta$ to create a protected version $x_{adv}=x_{ref}+\delta$. This perturbation is optimized to disrupt the diffusion synthesis process when conditioned on $x_{adv}$. Formally, we aim to find an optimal $\delta$ that maximizes the dissimilarity between the outputs generated using $x_{ref}$ and $x_{adv}$.
$\arg\max_{\delta}\ \mathcal{D}(\mathcal{VC}(x_{src},x_{ref},t),\ \mathcal{VC}(x_{src},x_{adv},t)),$ (4)
$\text{subject to}\quad\|\delta\|_{\infty}\leq\epsilon,$
where $\mathcal{D}(\cdot)$ measures the output discrepancy, $t$ is the diffusion timestep, and $\epsilon$ is the $l_{\infty}$-norm budget for the perturbation $\delta$. Our approach focuses on designing specific adversarial objectives that implicitly define $\mathcal{D}$ by exploiting intrinsic vulnerabilities within the diffusion mechanism itself.
4.2 Adversarial Identity Obfuscation
Opposite-Gender Embedding Centroid Guidance
Inspired by prior work showing dedicated speaker embeddings capture identity (Chen et al. 2024), directly manipulating these embeddings offers a direct approach, but its practical utility is hindered by poor transferability when attacking unknown or different encoder models.
Therefore, we explore leveraging speech Representation Learning (Chen et al. 2022b) to extract general acoustic representations that inherently capture speaker identity cues. Specifically, we select WavLM (Chen et al. 2022a) as our representation extractor, denoted as $\mathcal{R}(\cdot)$. By applying perturbations within this more general feature space, we aim for broader effectiveness against various models. As a baseline untargeted objective, we consider maximizing the embedding distance:
$\mathcal{L}_{ID}=1-Sim(\mathcal{R}(x_{adv}),\mathcal{R}(x_{ref})),$ (5)
where $Sim(x_{1},x_{2})$ represents the cosine similarity metric. While this untargeted objective effectively pushes $x_{adv}$ away from the reference identity representation domain, it lacks specific directionality in high-dimensional space. So we further incorporate a targeted component guided by psychoacoustic principles.
Psychoacoustic studies (Kreiman and Sidtis 2011; Lavner, Gath, and Rosenhouse 2000) indicate that significant perceptual differences in speaker identity often exist between genders, linked to specific acoustic cues like F0 and formant structures . Leveraging this, we assume that guiding the adversarial embedding towards the opposite gender one will likely create the strongest perceptual contrast, enhancing identity obfuscation.
Based on this insight, we propose an auditory-perception-guided adversarial perturbation. Specifically, we first randomly select a speaker from the dataset whose gender is opposite to that of $x_{ref}$. To establish a representative identity embedding, we compute the centroid $\mathcal{C}_{opp}$ by averaging the embeddings of all utterances in $\mathcal{X}$ within the WavLM feature space $\mathcal{R}$:
$\mathcal{C}_{opp}=\frac{1}{N}\sum_{x_{i}\in\mathcal{X}}\mathcal{R}(x_{i}),$ (6)
$\mathcal{L}_{ID}=-Sim(\mathcal{R}_{adv},\mathcal{R}_{ref}+\underbrace{Sim(\mathcal{R}_{adv},\mathcal{C}_{opp})}_{\text{Gender}},$ (7)
where $\mathcal{X}$ is the set of opposite gender utterances, $N$ represents length of $\mathcal{X}$, $\mathcal{R}_{i}$ is the embedding of $x_{i}$ in space $\mathcal{R}$.
Finally, minimizing the Eq. 7 directs the optimization process to simultaneously dissociate from the original speaker identity and converge towards a region selected based on psychoacoustic principles to maximize identity ambiguity.
Attention Context Divergence
Building upon the motivation to disrupt conditional guidance for identity obfuscation (Section 3.2), we introduce Attention Context Divergence. This strategy targets the attention mechanism to interfere with its use of contextual information from $x_{ref}$.
Commonly in diffusion-based VC, conditional information from $x_{ref}$ is integrated via Linear-attention layers (Katharopoulos et al. 2020). Within the U-Net, latent code representing the content $x_{src}^{t}$ are linearly projected to $Q$ matrix, whereas condition latents $x_{ref}^{t}$ are projected to $K$ and $V$. Linear-attention first computes a context matrix by aggregating values ($V$) weighted by keys ($K$):
$\mathcal{H}_{ctx}(x_{ref})=\text{Softmax}(\phi^{(l,t)}(x_{ref}^{t})W_{K}^{l})(\phi^{(l,t)}(x_{ref}^{t})W_{V}^{l})^{T},$
where $\mathcal{H}_{ctx}$ represents the context hidden state, $\phi^{(l,t)}(\cdot)$ is the deep features of the $l^{th}$ block in U-Net at timestep $t$, and $W_{K}^{l},W_{V}^{l}$ are the projection matrices. Then, the context interacts with queries ($Q$) to obtain the final attention output: $\mathcal{A}^{l}=(\mathcal{H}_{ctx})^{T}(W_{Q}^{l}\phi^{(l,t)}(x_{src}^{t}))$. This context representation provides a dynamic summary of the reference speaker’s stylistic features, weighted by their relevance to the current content queries during synthesis. By applying a softmax function, we obtain an explicit probability distribution over sequence positions which signifies the model’s information focus. Our strategy, therefore, is to maximize the Kullback-Leibler (KL) divergence (Kullback and Leibler 1951) between the context distribution derived from the original reference and the adversarial audio:
$\mathcal{L}_{ctx}$ $=D_{KL}(P_{ref}\parallel P_{adv}),$ (8)
$\text{where}\ P_{adv}$ $=\text{Softmax}(\mathcal{H}_{ctx}(x_{adv})),$
Maximizing this divergence forces the attention pattern to deviate from the original, thereby hindering accurate style alignment.
To enhance the impact on identity, we directs the adversarial pressure on the U-Net’s downsampling path. The rationale is that these earlier layers primarily process coarser, lower-frequency features (Wang et al. 2024) strongly associated with speaker timbre and identity. Therefore, focusing our proposed loss $\mathcal{L}_{ctx}$ on the attention layers within the U-Net’s downsampling path, the calculation of $P_{adv}$ can be restated as:
$P_{adv}=\text{Softmax}(\sum_{l}\mathcal{H}_{ctx}^{l}(x_{adv})),$ (9)
where the layer index $l$ iterates over the set Down ($l\in\text{Down}$), representing the U-Net’s downsampling blocks.
4.3 Perceptual Fidelity Degradation
Score Magnitude Amplification
To degrade perceptual fidelity by exploiting the sensitivity of the denoising trajectory, we introduce Score Magnitude Amplification. This design directly interferes with the score function $s_{\theta}$, which is estimated by the U-Net according to Eq. 3 and provides the essential drift term for the reverse SDE. We posit that the magnitude of $s_{\theta}$ relates to the strength of the drift guiding the noisy sample toward the target data manifold. Exploiting this connection, the SMA objective involves maximizing the magnitude of the score prediction $s_{\theta}$:
$\mathcal{L}_{score}=\mathbb{E}_{p_{t}(x),t\sim\mathcal{U}(1,T_{adv})}[\|s_{\theta}(x_{src}^{t},x_{adv}^{t},t)\|_{2}],$ (10)
where $p_{t}(x)$ is the distribution of noisy samples $x^{t}\sim q(x^{t}|x^{0})$, $\mathbb{E}[\cdot]$ calculates the average value, $T_{adv}$ stands for adversarial timesteps which will be discussed in Section 4.4. Optimizing the above formula introduces an erroneous drift strength. Consequently, the denoising trajectory is forcefully diverted, resulting in a collapse in perceptual quality.
Furthermore, the iterative nature of the diffusion process may amplify these induced trajectory deviations. Perturbations introduced at earlier timesteps can propagate through subsequent steps. This error cumulative effect thus enhances the efficacy of our adversarial strategy.
Noise-Guided Semantic Corruption
Following the motivation outlined in Section 3.2, we introduce a bidirectional semantic interference strategy. The core idea is twofold: (1) compel the features generated with $x_{adv}$ to diverge from those generated using the original reference $x_{ref}$, and (2) concurrently guide these adversarial features towards a ”semantic-free” state.
Specifically, consider a network layer $l$ within the frozen U-Net $\mathcal{U}_{\theta}$ and timestep $t$, we extract the original features $f^{(l,t)}=\mathcal{U}_{\theta}^{l}(x_{src},x_{ref},t)$ conditioned on $x_{ref}$, and $f^{(l,t)}_{adv}=\mathcal{U}_{\theta}^{l}(x_{src},x_{adv},t)$ corresponding to the adversarial version. Furthermore, to define a ”semantic-free” target, we leverage the U-Net’s activation modes to unstructured information. We extract features $f^{(l,t)}_{noise}=\mathcal{U}_{\theta}^{l}(x_{noise},x_{noise},t)$ by feeding Gaussian white noise $x_{noise}$ as both the source content and the reference condition. $f^{(l,t)}_{noise}$ can be considered to represent unstructured features and lack semantic information. The bidirectional objective aims to maximize the distance between $f^{(l,t)}_{adv}$ and $f^{(l,t)}$ while minimizing it with $f^{(l,t)}_{noise}$. This objective encourages the adversarial features to abandon the original semantic structure and move towards a state of incoherence which can be formalized as:
$\mathcal{L}_{sem}=1-\cos(f^{(l,t)}_{adv},f^{(l,t)})+\underbrace{\cos(f^{(l,t)}_{adv},f^{(l,t)}_{noise})}_{\text{Sem-free}},$ (11)
where we employ the cosine distance metric $\cos(\cdot)$ as it emphasizes the structural similarity between high-dimensional features rather than their absolute error. For enhanced impact on perceptual quality, the $\mathcal{L}_{sem}$ is strategically applied to layers within the U-Net’s upsampling path. These layers are critical for reconstructing the fine-grained acoustic details that govern output naturalness and perceived quality.
4.4 Joint Optimization of Defense Objectives
The final adversarial perturbation $\delta$ for VoiceCloak is optimized to simultaneously achieve our dual objectives (Section 4.2 and 4.3). The joint objectives of this comprehensive defense are formalized as follows
$\mathcal{L}_{total}=$ $(\mathcal{L}_{ID},\mathcal{L}_{ctx},\mathcal{L}_{score},\mathcal{L}_{sem})\Lambda^{T},$ (12)
$\delta$= $\ \arg\max_{\delta}\ \mathcal{L}_{total},$
where $\Lambda=(\lambda_{ID},\lambda_{ctx},\lambda_{score},\lambda_{sem})$ controls the weight factors that balance the relative importance of these defenses.
The efficacy of our perturbation optimization is influenced by the choice of diffusion timesteps $T_{adv}$ used for gradient computation. Informed by prior work (Yu et al. 2024) indicating early denoising steps primarily reconstruct low-frequency overall structural signal, we concentrate the optimization on these initial steps to maximize the disruption of fundamental integrity and reduce computational overhead.
Datasets Methods Defense Effectiveness Imperceptibility
[…]
LibriTTS Undefended - 76.49% - 3.96 - - - -
Random Noise $l_{\infty}$ 2.01 55.20% 0.31 3.72 16.00% 3.37 1.35 34.80
Attack-VC 2.29 36.20% 0.31 3.57 30.40% 2.31 3.71 5.29
VoicePrivacy 2.26 20.80% 0.30 3.60 26.80% 2.99 1.37 33.25
VoiceGuard 2.08 16.49% 0.29 3.63 43.45% 2.15 4.39 10.58
Ours 2.12 11.40% 0.27 2.36 71.40% 3.22 1.29 33.53
VCTK Undefended - 63.68% - 3.41 - - - -
Random Noise $l_{\infty}$ 1.68 58.00% 0.35 3.16 11.38% 3.25 1.38 34.10
Attack-VC 2.05 38.50% 0.33 2.82 26.20% 2.25 3.80 4.50
VoicePrivacy 1.88 30.28% 0.35 2.77 39.06% 2.87 1.53 31.87
VoiceGuard 1.87 31.42% 0.32 3.02 22.11% 2.05 4.57 10.00
Ours 1.93 19.74% 0.29 2.51 63.41% 3.09 1.33 32.41
Table 1: Comparison of defense effectiveness and adversarial imperceptibility with SOTA methods. Higher values are better for metrics marked with $\uparrow$, and vice versa for those marked with $\downarrow$. The best result is marked in BOLD.
5 Experiments
5.1 Experimental Setup
Datasets
Experiments are conducted on the LibriTTS (Zen et al. 2019) and VCTK (Yamagishi, Veaux, and MacDonald 2019) datasets. We selected a gender-balanced audio subsets (479 utterances from LibriTTS, 500 from VCTK) to generate adversarial reference speech set $\mathcal{D}_{x_{adv}}$.
Baseline Methods
We compare VoiceCloak against existing voice protection methods. We adopt the following methods for fair comparison: Attack-VC (Huang et al. 2021), VoicePrivacy (Chen et al. 2024), and VoiceGuard (Li et al. 2023). We also include a naive baseline: adding random Gaussian noise to $x_{ref}$.
Evaluation Metrics
For identity protection, we report the Automatic Speaker Verification (ASV) acceptance rate, where a lower rate for protected outputs signifies more effective obfuscation. We define a comprehensive Defense Success Rate (DSR) to measure the achievement of both our objectives. A defense is considered successful if the protected output both fails speaker verification and exhibits low perceptual quality ($\mathrm{DS}=s_{ASV}<\tau_{ASV}\land\mathrm{NISQA}(y_{adv})<\tau_{q}$), with thresholds $\tau_{ASV}=0.25$ and $\tau_{q}=3.0$. Additional metrics include Dynamic Time Warping (DTW) (Sakoe and Chiba 1978) and SSIM between $y$ and $y_{adv}$ spectrograms. Perturbation imperceptibility on $x_{adv}$ versus $x_{ref}$ is measured using PESQ (Rix et al. 2001), Mel-Cepstral Distortion (MCD) (Kubichek 1993) and SNR.
Implementation Details
We conduct our experiments mainly using DiffVC (Popov et al. 2022) as the target system. We set the number of optimization iterations to 50 and a step size $\alpha=4\times 10^{-5}$. And $\delta$ is constrained within an $l_{\infty}$-norm ball of $\epsilon=0.002$. We set the adversarial and inference timesteps respectively $T_{adv}=6$ and $T=100$. The loss function combined identity and quality objectives with weights $\Lambda=(1.0,4.5,10,0.85)$. All experiments are conducted on NVIDIA RTX 3090 GPU with a fixed random seed.
5.2 Comparison and Analysis
Comparison with Baselines
Figure 3: Mel spectrograms with $F_{0}$ pitch contours (green lines), and inferred intonation of the corresponding words. Arrows indicate perceived intonation shifts. (Intonation aligns with the ground truth, which is marked by green arrows, and diverges, which is marked by red arrows.)
As shown in Table 1, VoiceCloak demonstrates exceptional defense efficacy, significantly outperforming all baselines with a Defense Success Rate (DSR) of over 71% on LibriTTS and 63% on VCTK. This high DSR reflects success in both of our defense objectives: identity obfuscation is achieved by drastically reducing speaker verification acceptance rates to  11%, while perceptual quality degradation is confirmed by low NISQA scores indicating unacceptable audio generation quality. This dual ability to effectively cripple both identity mimicry and audio usability distinguishes VoiceCloak from defenses that may struggle with diffusion models or focus primarily on one objective. As expected, the naive Random Noise baseline confirms that unstructured noise provides very limited protection. Regarding imperceptibility, VoiceCloak performs comparably to baselines. This supports our strategy of targeting intrinsic vulnerabilities, rather than simply increasing perturbation magnitude.
Figure 3 further visualizes disruption result. Applying VoiceCloak protection results markedly different from the undefended one. The $F_{0}$ curve is notably degraded, appearing blurred and unpredictable, accompanied by highly inconsistent intonation changes.
Protecting Commercial Systems
Figure 4: Protecting commercial speaker verification APIs (Iflytek, Azure) from spoofing attacks (lower are better).
We also evaluated the effectiveness in protecting commercial speaker verification (SV) APIs (Iflytek, Azure) to simulate real-world anti-spoofing scenarios. Successful protection aims to minimize the similarity score returned by the API. Figure 4 demonstrated VoiceCloak’s superior ability to decouple the protected audio from the original speaker’s identity.
User Study
Figure 5: User perceptual study results. (a) Timbre Dissimilarity Preference. (b) Corresponding results for perceived Naturalness Disruption.
We conducted a user study with 50 participants to assess perceptual impact. Listeners performed comparisons on two criteria: Timbre Dissimilarity and Naturalness Disruption and we aggregated results in Figure 5, where ”Neutral” indicates no perceived difference. Participants consistently rated VoiceCloak’s outputs as having both greater timbre dissimilarity and more severe naturalness disruption, confirming its human-perceived effectiveness.
Settings Defense Effectiveness
[…]
- - 1.96 46.82% 0.30 3.66 22.58%
$\checkmark$ - 2.16 8.57% 0.30 3.57 27.74%
w/o Gender - 2.25 19.92% 0.30 3.60 14.40%
- $\checkmark$ 2.31 16.20% 0.27 2.96 62.57%
$\checkmark$ $\checkmark$ 2.13 11.00% 0.27 2.85 69.20%
Table 2: Ablation study on the contribution of different settings for Identity Obfuscation. $\checkmark$ indicates the setting is used, ”w/o” denotes the exclusion of the specified term.
5.3 Ablation Study
Effectiveness of Adversarial Identity Obfuscation
Settings Defense Effectiveness
[…]
- - 1.99 45.00% 0.31 3.09 20.20%
$\checkmark$ - 2.42 31.80% 0.29 2.68 41.20%
- $\checkmark$ 2.23 23.00% 0.27 2.44 60.60%
- w/o Sem-free 2.28 26.36% 0.29 3.30 26.80%
$\checkmark$ $\checkmark$ 2.22 23.60% 0.27 2.10 57.80%
Table 3: Ablation study for Perceptual Fidelity Degradation. Checkmark ($\checkmark$) indicates the setting is used,”w/o” denotes the exclusion of the specified loss term.
We ablate the loss designed for Adversarial Identity Obfuscation, $\mathcal{L}_{ID}$ and $\mathcal{L}_{ctx}$, in Table 2. The results show that $\mathcal{L}_{ID}$ alone effectively reduces the ASV acceptance rate. This confirms that directly manipulating the representation learning embedding space effectively disrupts recognizable identity features. Removing the opposite-gender guidance from $\mathcal{L}_{ID}$ worsens ASV, confirming our psychoacoustically-motivated strategy provides effective direction for identity disruption. Separately, the context divergence loss also contributes significantly, lowering both ASV and NISQA by interfering with the attention mechanism’s condition injection.
Effectiveness of Perceptual Fidelity Degradation
Table 3 ablates the perceptual quality degradation losses. The results show that $\mathcal{L}_{score}$ alone is effective, degrading quality (lower NISQA, higher DTW) by forcing the denoising trajectory away from high-fidelity regions. The semantic corruption loss, $\mathcal{L}_{score}$, demonstrates an even stronger individual impact by directly corrupting internal U-Net features, thereby impairing the model’s reconstruction of coherent, natural-sounding speech details. Critically, the necessity of guiding semantic features towards an incoherent state is evident from the ”w/o Sem-free” variant.
5.4 Transferability
Target Models Defense Effectiveness
[…]
DiffVC 2.12 11.40% 0.27 2.36 71.40%
DDDM-VC 1.67 16.80% 0.36 2.79 54.89%
DuTa-VC 2.41 13.77% 0.27 2.14 73.92%
Table 4: Transferability of the proposed defense method against unseen target models.
To demonstrate applicability, we extended our experiments to include two additional open-source diffusion-based VC models: DuTa-VC (Wang et al. 2023) and DDDM-VC (Choi, Lee, and Lee 2024). As shown in Table 4, VoiceCloak demonstrate favorable transferability to different models, achieving an average DSR of 66.7%. We attribute this transferability to: Targeting Common Vulnerabilities, as our method exploits fundamental mechanisms (e.g. attention, score prediction) that are shared across diffusion VCs.
5.5 Robustness
Figure 6: Resilience of VoiceCloak under more advanced robust scenario.
We evaluated the robustness of our method against four common distortions, with results presented in Figure 6. The results show consistent defense effectiveness, indicating resilience to real-world transformations and significantly outperforming the undefended baseline. Further details and analysis are available in Appendix.
5.6 Efficiency Discussion
Time (s) Mem. (GiB)
Ours 148.66 8.97
Table 5: Computational overhead analysis. ”Time”: average time to generate one sample. ”Mem.”: peak GPU usage.
As shown in Table 5, the protection process is a one-time, offline operation, and the required GPU memory is within the capacity of consumer-grade hardware, making VoiceCloak a practical defense method.
6 Conclusion
This paper introduced VoiceCloak, a comprehensive defense against unauthorized diffusion-based voice cloning (VC). Our framework achieves superior defense effectiveness by exploiting targeted intrinsic vulnerabilities within the diffusion process. Through strategies designed to disrupt attention-based conditional guidance, steer the denoising trajectory, and corrupt internal semantic representations, VoiceCloak effectively undermines the synthesis process. Extensive experiments validate its efficacy, demonstrating success in simultaneously obfuscating speaker identity and degrading audio quality to mitigate the threats of voice mimicry Extensive experiments validate its ability to significantly hinder voice cloning by simultaneously disrupting identity and degrading audio quality, thereby mitigating the threats of high-quality voice mimicry